### v1

# On the strong DR/DZ equivalence conjecture

Abstract.
We establish the Miura equivalence of two integrable systems associated to a semi-simple cohomological field theory: the double ramification hierarchy of Buryak and the Dubrovin-Zhang hierarchy. This equivalence was conjectured by Buryak and further refined by Buryak, Dubrovin, Guéré, and Rossi.
1. Introduction
1.1. Cohomological field theories
Let $V=\langle e_{1},\dots,e_{N}\rangle$ be a finite dimensional vector space equipped with a symmetric nondegenerate bilinear form $\eta$. A cohomological field theory (CohFT) [23] is a system of linear maps
(1.1) $c_{g,n}\colon V^{\otimes n}\to H^{*}({\overline{\mathcal{M}}}_{g,n}),\quad 2g-2+n>0,$
such that
  1. (1)
    The maps $c_{g,n}$ are equivariant with respect to the $S_{n}$-action permuting the factors of $V$ in $V^{\otimes n}$ and the labels of the marked points in ¯${\overline{\mathcal{M}}}_{g,n}$.
  2. (2)
    $\pi^{*}c_{g,n}(\otimes_{i=1}^{n}e_{\alpha_{i}})=c_{g,n+1}(\otimes_{i=1}^{n}e_{\alpha_{i}}\otimes e_{1})$, where $\pi\colon{\overline{\mathcal{M}}}_{g,n+1}\to{\overline{\mathcal{M}}}_{g,n}$ is the map that forgets the last marked point. Moreover, $c_{0,3}(e_{\alpha}\otimes e_{\beta}\otimes e_{1})=\eta(e_{\alpha}\otimes e_{\beta})=:\eta_{\alpha\beta}$.
  3. (3)
    $\mathrm{gl}^{*}c_{g,n}(\otimes_{i=1}^{n}e_{\alpha_{i}})=\eta^{\mu\nu}c_{g_{1},n_{1}+1}(\otimes_{i\in I_{1}}e_{\alpha_{i}}\otimes e_{\mu})\otimes c_{g_{2},n_{2}+1}(\otimes_{j\in I_{2}}e_{\alpha_{j}}\otimes e_{\nu})$, where $I_{1}\sqcup I_{2}=\{1,\dots,n_{1}+n_{2}\}$, $|I_{1}|=n_{1}$, $|I_{2}|=n_{2}$, $g_{1}+g_{2}=g$, and $\mathrm{gl}\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+1}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+1}\to{\overline{\mathcal{M}}}_{g,n}$ is the corresponding gluing map.
  4. (4)
    $\mathrm{gl}^{*}c_{g,n}(\otimes_{i=1}^{n}e_{\alpha_{i}})=\eta^{\mu\nu}c_{g-1,n+2}(\otimes_{i=1}^{n}e_{\alpha_{i}}\otimes e_{\mu}\otimes e_{\nu})$, where $\mathrm{gl}\colon{\overline{\mathcal{M}}}_{g-1,n+2}\to{\overline{\mathcal{M}}}_{g,n}$ is the corresponding gluing map.
CohFTs capture the algebraic properties of enumerative problems in different contexts, among which Gromov-Witten theory. They are fully classified in the case $c_{0,3}$ defines a structure of a semi-simple Frobenius algebra on $V$ in [26]. In particular, under this assumption they consist of tautological classes.
One associates to a CohFT its potential
(1.2) $F\coloneqq\sum\frac{\varepsilon^{2g}}{n!}\left(\int_{{\overline{\mathcal{M}}}_{g,n}}c_{g,n}(\otimes_{i=1}^{n}e_{\alpha_{i}})\prod_{i=1}^{n}\psi_{i}^{d_{i}}\right)\prod_{i=1}^{n}t^{\alpha_{i},d_{i}}$
and a partition function $Z\coloneqq\exp(\epsilon^{-2}F)$.
1.2. Dubrovin-Zhang hierarchy
One can associate to a semi-simple CohFT a Hamiltonian integrable hierarchy of hydrodynamic type. This construction was initially developed by Dubrovin and Zhang [18], and subsequently by Liu, Wang, and Zhang [24], in the framework of calibrated Frobenius manifolds. In the context of CohFTs, calibration refers to an extra homogeneity property.
A more general approach can be applied to any semi-simple CohFT and it is developed in [10, 11]. In a nutshell, one can define a new system of coordinates
(1.3) $w^{\alpha,d}\coloneqq\partial_{x}^{d}\eta^{\alpha\beta}\frac{\partial^{2}F}{\partial t^{\beta,0}t^{1,0}}\big{|}_{t^{1,0}\to t^{1,0}+x}.$
The double derivatives of $F$ are proved to be differential polynomials in these new coordinates,
(1.4) $\frac{\partial^{2}F}{\partial t^{\alpha,p}\partial t^{\beta,q}}=\Omega_{\alpha,p;\beta,q}(\{w^{\gamma,d}\}),$
and the following evolutionary integrable system possesses a polynomial Hamiltonian structure of hydrodynamic type with a Poisson bracket $K^{\mathrm{DZ}}=\eta^{\alpha\beta}{\partial}_{x}+O(\epsilon^{2})$ and a tau structure:
(1.5) $\frac{\partial w^{\alpha}}{\partial t^{\beta,q}}=\partial_{x}\eta^{\alpha\gamma}\Omega_{\gamma,0;\beta,q}.$
This system is described explicitly in a number of examples: for the trivial CohFT it is the KdV hierarchy [27, 22], see also [2], and for the Witten $r$-spin CohFT it is the Gelfand-Dickey hierarchy [28, 19].
1.3. Buryak’s double ramification hierarchy
Buryak constructed a different integrable system associated to a CohFT in [5]. Let $u^{1},\ldots,u^{N}$ be formal dependent variables, and let $u^{\alpha,d}\coloneqq{\partial}_{x}^{d}u^{\alpha}$. Define differential polynomials $P^{\alpha}_{\beta,d}$ by
[…]
where $I^{g,d}_{\gamma,\beta,\alpha_{1},\dots,\alpha_{n}}$ is a polynomial of degree $2g$ in the variables $a_{1},\dots,a_{n}$ given by
(1.7) $I^{g,d}_{\gamma,\beta,\alpha_{1},\dots,\alpha_{n}}\coloneqq\left(\int_{{\overline{\mathcal{M}}}_{g,n+2}}\lambda_{g}{\mathrm{DR}_{g}\bigg{(}0,-\sum_{j=1}^{n}a_{j},a_{1},\ldots,a_{n}\bigg{)}}\psi_{1}^{d}c_{g,n+2}\Big{(}e_{\beta}\otimes e_{\gamma}\otimes\bigotimes_{j=1}^{n}e_{\alpha_{j}}\Big{)}\right).$
Buryak’s DR hierarchy [5] is the following integrable system of evolutionary PDEs:
(1.8) $\frac{{\partial}u^{\alpha}}{{\partial}t^{\beta,d}}={\partial}_{x}P^{\alpha}_{\beta,d}.$
It possesses the Hamiltonian structure given by $\eta^{\alpha\beta}{\partial}_{x}$ and a tau structure. Note, however, that $u^{\alpha}$ are not in general the normal coordinates for the tau structure.
The further properties of this integrable system and its generalizations are studied in a number of papers, see [13, 6, 7, 15]. The importance of the Buryak’s approach to integrable systems associated to semi-simple CohFTs is in particular justified by that fact that it admits a quantization with respect to its Poisson structure, see [12, 7].
1.4. Normal Miura equivalence
In general the Miura group acts on (Hamiltonian) integrable systems by the diffeomorphic changes of dependent variables, see e.g. [18]. More refined are the so-called normal Miura transformations that transform a tau-symmetric Hamiltonian hierarchy written in normal coordinates to a tau-symmetric Hamiltonian hierarchy written in normal coordinates.
Let $\tilde{u}^{\alpha}$ be the normal coordinates for Buryak’s double ramification hierarchy. The normal coordinates for the Dubrovin-Zhang hierarchy are $w^{\alpha}$. The following conjecture is proposed in [6, Conjecture 7.3]:
Conjecture 1.1 (Strong DR/DZ equivalence conjecture).
For any semi-simple cohomological field theory there exists a differential polynomial $\mathcal{P}$ in coordinates $w^{\alpha}$ (and it has a very explicit description that we omit here) such that the normal Miura transformation defined as
(1.9) $w^{\alpha}\mapsto\tilde{u}^{\alpha}(w)=w^{\alpha}+\eta^{\alpha\gamma}{\partial}_{x}\{\mathcal{P},\bar{h}^{\mathrm{DZ}}_{\mu,0}\}_{K^{DZ}}$
maps the Dubrovin-Zhang hierarchy to Buryak’s double ramification hierarchy written in the normal coordinates.
In the statement of the conjecture $\bar{h}^{\mathrm{DZ}}_{\mu,0}$ is the Hamiltonian of the time flow $\partial_{t^{\mu,0}}$ for the Dubrovin-Zhang hierarchy; {,}$\{\bullet,\bullet\}_{K^{DZ}}$ denotes the Poisson bracket of the Dubrovin-Zhang hierarchy. This conjecture refines an earlier conjecture of Buryak [5] on Miura equivalence of the Dubrovin-Zhang and double ramification hierarchies.
The main application of our work is the following theorem:
Theorem 1.2.
Conjecture 1.1 holds.
1.5. Generalizations and approach to the proof
There is a number of generalizations of the strong DR/DZ equivalence conjecture for semi-simple CohFTs above. In particular, the condition of semi-simplicity can be dropped (but then one has to prove that the Dubrovin-Zhang integrable system is given by differential polynomials in the dependent variables), and similar constructions can be applied to the so-called partial CohFTs and F-CohFTs (in the latter case we get just a system of conservation laws rather than a Hamiltonian system). These constructions are available in [1, 14, 16], and the corresponding conjectures there state that
  1. (1)
    On the Dubrovin-Zhang side, the corresponding integrable system is represented by differential polynomials.
  2. (2)
    The Dubrovin-Zhang integrable system is normal Miura equivalent to the double ramification system.
We refer to [16, Section 4] for the precise statement and further details. In [16], following and generalizing earlier results in [6, 8], these conjectures are reduced to a system of conjectural tautological relations in the moduli space of curves known as a generalization of the $A=B$ relation. The initial $A=B$ relation itself was proposed in [8] and designed to imply Conjecture 1.1. It is convenient to let the term “$A=B$ relation” refer to the whole system of (generalized) $A=B$ relations.
In this paper we prove the $A=B$ relation as well as all its generalization from [16] in the Gorenstein quotient. In other words we prove the relations after intersection with tautological classes of the complementary dimension (Theorem 2.3). As a consequence, our result implies the entire system of conjectures for not-necessary semi-simple CohFTs, as well as partial CohFTs and F-CohFTs under the assumption that the classes involved in these structures are tautological. Since this is always the case for semi-simple CohFTs, as a corollary Theorem 2.3 implies Conjecture 1.1 (Theorem 1.2). We remark that Theorem 2.3 is the main result of this paper, and we moreover refer further to [16, Section 4] for a list of corollaries descending from it.
1.6. Structure of the paper
This paper is by no means self-closed as it is based on an enormous amount of prior work; we expect the reader to be familiar with the theory of (infinite-dimensional) integrable systems of evolutionary type along the lines of [18] and [6] as well as the intersection theory of the moduli spaces of curves and the structure of its tautological ring, along the lines of [20] and [17]. Moreover, it might be instructive to follow the big steps in the literature where the conjecture of Buryak and its generalizations were further developed, which we summarise in the following.
  • Buryak’s double ramification hierachy and the original conjecture are formulated in [5]; the conjecture is proved there for the trivial CohFT;
  • The conjecture is refined to be an explicitly presented normal Miura transformation in [6]; the strong DR/DZ equivalence is proved there for the trivial CohFT;
  • The refined conjecture is reduced to a system of relations in the tautological ring of the moduli space of curves in [8];
  • The tautological relations is further generalized to a form that we study here and is connected to a number of more general settings explained in [16];
  • The relations expressions are slightly simplified in [3], in line with a twin system of relations that involve the so-called $\Omega$-classes;
  • Some new combinatorial insights, that in particular allowed to prove the twin system of tautological relations, are obtained in [4];
  • Finally, here we prove the desired relations in the Gorenstein quotient.
Starting from this point we only talk about classes in the tautological ring of the moduli space of curves. In Section 2 we introduce the conjectural tautological relations, the generalized $A=B$ relations, in a bit more compact form than they were presented before, and recall some basic facts about their intersections with $\psi$-classes that follow from the strong DR/DZ equivalence for the trivial CohFT.
In Section 3 we introduce a new system of conjectural tautological relations, the so-called master relations, which are strongly inspired by our work on the twin system of tautological relations and localization techniques. We also make some steps towards establishing their vanishing in the Gorenstein quotient.
Finally, in Section 4 we use a combinatorial relation between the generalized $A=B$ relations and the master relations in two different directions. In one direction, we show that the generalized $A=B$ relations are equivalent to the master relations, and in particular this equivalence holds in the Gorenstein quotient. And in the other direction, we transfer the properties of $A=B$ cycles implied by the DR/DZ correspondence for the trivial CohFT to complete the proof that the master relations vanish in the Gorenstein quotient. This proves that the generalized $A=B$ relations hold in the Gorenstein quotient.
1.7. Acknowledgments
X. B. and S. S. were supported by the Netherlands Organization for Scientific Research. D. L. is supported by the University of Trieste, by the INFN under the national project MMNLP (APINE) section of Trieste, by the INdAM group GNSAGA, and by the PRIN project 2022 “Geometry of algebraic structures: moduli, invariants, deformations”. The authors thank A. Buryak and P. Rossi for useful discussions and collaboration on closely related topics.
2. The $A=B$ relation in the Gorenstein quotient
2.1. Basic notation for trees
Let $\mathrm{SRT}_{g,n,m}$ be the set of stable rooted trees of total genus $g$, with $n$ regular legs $\sigma_{1},\dots,\sigma_{n}$ and $m$ extra legs $\sigma_{n+1},\dots,\sigma_{n+m}$, which we refer to as “frozen” legs and must always be attached to the root vertex. For a $T\in\mathrm{SRT}_{g,n,m}$ we use the following notation:
  • $H(T)$ is the set of half-edges of $T$.
  • $L(T),L_{r}(T),L_{f}(T)\subset H(T)$ are the sets of all, regular, and frozen legs of $T$, respectively. $L(T)=L_{r}(T)\sqcup L_{f}(T)$.
  • $H_{e}(T)\coloneqq H(T)\setminus L(T)$.
  • $\iota\colon H_{e}(T)\to H_{e}(T)$ is the involution that interchanges the half-edges that form an edge.
  • $E(T)$ is the set of edges of $T$, $E\cong H_{e}(T)/\iota$.
  • $H_{+}(T)\subset H(T)$ is the set of the so-called “positive” half-edges that consists of all regular legs of $T$ and of half-edges in $H(T)\setminus L(T)$ directed away from the root at the vertices where they are attached, $H_{+}(T)\cong E(T)\cup L_{r}(T)$;
  • $H_{-}(T)\subset H(T)$ is the set of the so-called “negative” half-edges that consists of all frozen legs of $T$ and of half-edges in $H(T)\setminus L(T)$ directed towards the root at the vertices where they are attached, $H_{-}(T)\cong E(T)\cup L_{f}(T)$;
  • $V(T),V_{nr}(T)$ are the sets of vertices and non-root vertices of $T$.
  • $v_{r}\in V(T)$ is the root vertex of $T$; $V(T)=\{v_{r}(T)\}\sqcup V_{nr}(T)$.
  • For a $v\in V(T)$, $H(v),H_{+}(v),H_{-}(v)$ are all, positive, and negative half-edges attached to $v$, respectively. Obviously, $|H_{-}(v_{r})|=m$ and for any $v\in V_{nr}(T)$ we have $|H_{-}(v)|=1$.
  • For a $v\in V(T)$ let $g(v)\in\mathbb{Z}_{\geq 0}$ be the genus assigned to $v$. The stability condition means that
    $\chi(v)\coloneqq 2g(v)-2+|H(v)|>0.$
    The genus condition reads
    $\sum_{v\in V(T)}g(v)=g.$
  • We say that a vertex or a (half-)edge $x$ is a descendant of a vertex or a (half-)edge $y$ if $y$ is on the unique path connecting $x$ to $v_{r}$.
  • For an $h\in H_{+}(T)$ let $DL(h)$ be the set of all legs that are descendants to $h$, including $h$ itself. Note that $DL(h)\subseteq L_{r}(T)$ for any $h\in H_{+}(T)$ and $DL(l)=\{l\}$ for $l\in L_{r}(T)$.
  • For an $h\in H_{+}(T)$ let $DH(h)$ be the set of all positive half-edges that are descendants to $h$, excluding $h$. For instance, for $l\in L_{r}(T)$ we have $DH(l)=\emptyset$, and for $h\in H_{+}(T)\setminus L_{r}(T)$ we have $DH(h)\supseteq DL(h)$.
  • For an $e\in E(T)$ let $DL(e)$ be the set of all legs that are descendants to $e$. Note that $DL(e)\subseteq L_{r}(T)$ for any $e\in E(T)$.
  • For an $v\in V(T)$ let $DL(v)$ be the set of all regular legs that are descendants to $v$. In particular, $DL(v_{r})=L_{r}(T)$.
  • For a $v\in V(T)$ let $DV(v)\subset V(T)$ be the subset of all vertices that are descendants of $v$, including $v$ itself. For instance, $DV(v_{r})=V(T)$. Let
    $D\chi(v)\coloneqq\sum_{v^{\prime}\in DV(v)}\chi(v^{\prime}).$
Consider the polynomial ring $Q\coloneqq\mathbb{Q}[a_{1},\dots,a_{n}]$ and define $a\colon H_{+}(T)\to Q$, $a\colon E(T)\to Q$, and $a\colon V(T)\to Q$ (abusing notation we use the same symbol $a$ for all these maps) by
$a(\sigma_{i})$ $\coloneqq a_{i},$ $i=1,\dots,n;$ $a(h)$ $\coloneqq\textstyle\sum_{l\in DL(h)}a(l),$ $h\in H_{+}(T);$
$a(e)$ $\coloneqq\textstyle\sum_{l\in DL(e)}a(l),$ $e\in E(T);$ $a(v)$ $\coloneqq\textstyle\sum_{l\in DL(v)}a(l),$ $v\in V(T).$
2.2. Leveled stable rooted trees
We enhance the structure of a stable rooted tree to what we call a degree-labeled stable rooted tree (of genus $g$, with $n$ regular and $m$ frozen legs). To this end we take a stable rooted tree $T\in\mathrm{SRT}_{g,n,m}$ and assign to each $v\in V(T)$ an extra degree label $p(v)\in\mathbb{Z}_{\geq 0}$ such that $p(v)\leq 3g(v)-3+|H(v)|$. Let $\mathcal{P}(T)$ denote the set of degree label functions on $T$. A degree-labeled stable rooted tree is a pair $(T,p)$, where $T\in\mathrm{SRT}_{g,n,m}$ and $p\in\mathcal{P}$.
Our next goal is to assign to a degree-labeled stable rooted tree $(T,p)$ a so-called admissible level function. A function $\ell\colon V(T)\to\mathbb{Z}_{\geq 0}$ is called an admissible level function if the following conditions are satisfied:
  • The value of $\ell$ on the root vertex is zero ($\ell(v_{r})=0$).
  • If $v^{\prime}\in DV(v)$ and $v^{\prime}\not=v$, then $\ell(v^{\prime})>\ell(v)$.
  • There are no empty levels, that is, for any $0\leq i\leq\max\ell(V(T))$ the set $\ell^{-1}(i)$ is non-empty.
  • For every $10\leq i\leq\max\ell(V(T))-$ 1 we have inequality
    |$|\{v\in V(T)\,|\,\ell(v)\leq i\}|-1+\sum_{\begin{subarray}{c}v\in V(T)\\ \ell(v)\leq i\end{subarray}}p(v)\leq\sum_{\begin{subarray}{c}v\in V(T)\\ \ell(v)\leq i\end{subarray}}2g(v)-2+m.$
    For instance, if $T$ has more than one vertex, then $p(v_{r})\leq 2g(v_{r})-2+m$.
Let $\mathcal{L}(T,p)$ denote the set of admissible level functions on $(T,p\in\mathcal{P}(T))$.
The set $\mathrm{LDLSRT}_{g,n,m}$ of leveled degree-labeled stable rooted trees consists of triples $(T,p,\ell)$, where $T\in\mathrm{SRT}_{g,n,m}$, $p\in\mathcal{P}(T)$, and $\ell\in\mathcal{L}(T,p)$.
2.3. Definition of the $B$ class
The goal of this section is to define the so-called $B$ class in the tautological ring of ¯${\overline{\mathcal{M}}}_{g,n+m}$.
Let $(T,p,\ell)\in\mathrm{LDLSRT}_{g,n,m}$. Assign to each $v\in V(T)$ the moduli space of curves ¯${\overline{\mathcal{M}}}_{g(v),|H(v)|}$, where the first $|H_{+}(v)|$ marked points correspond to the positive half-edges attached to $v$ and ordered in an arbitrary but fixed way and the the last $|H_{-}(v)|$ marked points correspond to the negative half-edges attached to $v$, also ordered in some arbitrary but fixed way. Consider the class
$\Psi(v)\coloneqq\left(\prod_{i=1}^{|H_{+}(v)|}\frac{1}{1-a(h_{i})\psi_{i}}\right)_{p(v)}\in R^{p}({\overline{\mathcal{M}}}_{g(v),|H(v)|})\otimes_{\mathbb{Q}}Q$
For each $(g,n.m)$ such that $02g-2+n+m>$ 0 define the class
$B^{m}_{g,n}\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$
as
(2.1) $B^{m}_{g,n}\coloneqq\sum_{(T,p,\ell)\in\mathrm{SRT}(g,n,m)}(-1)^{\max\ell(V(T))}\biggl{(}\prod_{e\in E(T)}a(e)\biggr{)}(b_{T})_{*}\bigotimes_{v\in V(T)}\Psi(v)$
Here $(b_{T})_{*}$ is the boundary pushforward map that acts from $\bigotimes_{v\in V(T)}R^{*}({\overline{\mathcal{M}}}_{g(v),|H(v)|})\otimes_{\mathbb{Q}}Q$ to $R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$. The class $B^{m}_{g,n}$ has the feature that its component in $R^{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$, $d=0,\dots,3g-3+n+m$.
2.4. The conjectural $A=B$ relation revisited
Consider the moduli space
(2.2) ¯$\overline{\mathcal{M}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})$
of rubber stable maps to $(\mathbb{P}^{1},0,\infty)$. Let $s\colon{\overline{\mathcal{M}}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})\to{\overline{\mathcal{M}}}_{g,n+1}$ be the projection to the source curve, and $\lambda_{g}$ the lift of the lambda class with respect to this projection. Let $t\colon{\overline{\mathcal{M}}}_{g}^{\sim}(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})\to LM_{2g-1+n}$ be the projection to the target curve, where $LM_{2g-1+n}$ denotes the Losev-Manin space with $2g-1+n$ marked points. Let $\tilde{\psi}_{0}$ be the pull-back by $t$ of the $\psi$-class at the point 00 in Losev-Manin space. Define
(2.3) $A^{1}_{g,n}\coloneqq s_{*}\left(\frac{\lambda_{g}}{1-\tilde{\psi}_{0}}\left[{\overline{\mathcal{M}}}_{g}^{\sim}\left(\mathbb{P}^{1},a_{1},\dots,a_{n},-\sum\nolimits_{i=1}^{n}a_{i}\right)\right]^{\mathrm{vir}}\right)\in R^{*}({\overline{\mathcal{M}}}_{g,n+1})\otimes_{\mathbb{Q}}Q.$
Note that $(A^{1}_{g,n})_{<2g}=0$ and $(A^{1}_{g,n})_{2g}=\lambda_{g}\mathrm{DR}_{g}(a_{1},\dots,a_{n},-\sum_{i=1}^{n}a_{i})$. Note also that for any $d\geq 2g$ the class $(A^{1}_{g,n})_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
Conjecture 2.1.
The following three statements hold:
  1. (1)
    For any $(g,n,m)$ such that $g\geq 0$, $n\geq 1$, $m\geq 2$, we have
    (2.4) $\deg B^{m}_{g,n}\leq 2g-2+m.$
  2. (2)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-1+n>$ 0, we have
    (2.5) $\deg(B^{1}_{g,n}-A^{1}_{g,n})\leq 2g-1.$
  3. (3)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-2+n>$ 0, we have
    (2.6) $\deg(B^{0}_{g,n}-\frac{1}{\sum_{i=1}^{n}a_{i}}\pi_{*}A^{1}_{g,n})\leq 2g-2,$
    where $\pi\colon{\overline{\mathcal{M}}}_{g,n+1}\to{\overline{\mathcal{M}}}_{g,n}$ is the map that forgets the last marked point.
In all three statement $\deg$ refers either to cohomological degree or to the homogenenous degree in $a_{1},\dots,a_{n}$. We use the convention that a polynomial of negative degree is the zero polynomial. Note that the third statement follows from the second statement. We have the following lemma:
Lemma 2.2.
Conjecture 2.1 is equivalent to [16, Conjectures 1, 2, and 3].
Proof.
We have to show that our definition of the class $A^{1}_{g,n}$ coincides with the definition given in [16] (see also [3, Section 2.4]). To this end, we perform the intersection of the class $\tilde{\psi}_{0}$ with the moduli space of rubber maps using the technics of [17], this yields exactly the same trees in $\mathrm{SRT}_{g,n,1}$ decorated by double ramification cycles as in the original definition of $A^{1}_{g,n}$. The only difference is the combinatorial coefficient that we have to compare.
In the case of the original definition, the combinatorial coefficient of a tree $T\in\mathrm{SRT}_{g,n,1}$ is equal to
(2.7) $C_{1}(T)=\prod_{v\in V(T)}\frac{\chi(v)}{D\chi(v)}.$
For the computation of (2.3) we consider a set rigorous level structures $\mathcal{R}(T)$, that is, the set of one-to-one maps $f\colon V(T)\to\{1,\dots,|V(T)|\}$ such that $f(v_{1})>f(v_{2})$ if $v_{1}\in DV(v_{2})$. In particular, $f(v_{r})=1$. Then the coefficient implied by (2.3) is equal to
(2.8) $C_{2}(T)=\sum_{f\in\mathcal{R}(T)}\prod_{i=1}^{|V(T)|}\frac{\chi(f^{-1}(i))}{\chi(f^{-1}(i))+\cdots+\chi(f^{-1}(|V(T)|))}.$
Let us show that $C_{1}(T)=C_{2}(T)$. To this end, observe that both the definition of $C_{1}$ and $C_{2}$ can be applied to forests, moreover,
(2.9) $C_{1}(T_{1}\sqcup\cdots T_{k})=\prod_{i=1}^{k}C_{1}(T_{i})\quad\text{and}\quad C_{2}(T_{1}\sqcup\cdots T_{k})=\prod_{i=1}^{k}C_{2}(T_{i}).$
Moreover, the factor in $C_{1}(T)$ corresponding to $v_{r}$ coincides with the factor for $v_{r}$ in each summand in $C_{2}(T)$. Thus we can remove the root vertex and use the factorization for the forests above to prove $C_{1}(T)=C_{2}(T)$ by induction on the size of the tree. ∎
Various simplifications and partial results towards Conjecture 2.1 are available in [3, 9, 16, 21, 25].
2.5. The $A=B$ identity in the Gorenstein quotient
The main statement that we prove in this paper is the following theorem:
Theorem 2.3 (main theorem).
Conjecture 2.1 holds in the Gorenstein quotient. Namely, the following three statements hold:
  1. (1)
    For any $(g,n,m)$ such that $g\geq 0$, $n\geq 1$, $m\geq 2$, and for any $\alpha\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ we have
    (2.10) $\deg\left(\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha B^{m}_{g,n}\right)\leq 2g-2+m.$
  2. (2)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-1+n>$ 0, and for any $\alpha\in R^{*}({\overline{\mathcal{M}}}_{g,n+1})$ we have
    (2.11) $\deg\left(\int_{{\overline{\mathcal{M}}}_{g,n+1}}\alpha(B^{1}_{g,n}-A^{1}_{g,n})\right)\leq 2g-1.$
  3. (3)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-2+n>$ 0, and for any $\alpha\in R^{*}({\overline{\mathcal{M}}}_{g,n})$ we have
    (2.12) $\deg\left(\int_{{\overline{\mathcal{M}}}_{g,n}}\alpha(B^{0}_{g,n}-\frac{1}{\sum_{i=1}^{n}a_{i}}\pi_{*}A^{1}_{g,n})\right)\leq 2g-2.$
Remark 2.4.
The third statement is a corollary of the second one that one can obtain by the projection formula.
As a corollary, the first statement of this theorem implies the polynomiality of the system of Dubrovin-Zhang conservation laws associated to an arbitrary F-CohFT whose classes belong to the tautological ring. The second statement implies that the latter hierarchy is Miura equivalent, in a very explicit way, to the DR hierarchy of conservation laws associated to the same F-CohFT, also under assumption that its classes belong to the tautological ring. Finally, the third statement implies the strong DR/DZ correspondence for (partial) CohFTs, also under assumption that the classes of (partial) CohFT belong to the tautological ring. See [16, Section 4] and [6, 8].
One important lemma that allows to establish the strong DR/DZ correspondence for any semi-simple cohomological field theory is a direct corollary of the classification proved in [26]:
Lemma 2.5 (Corollary of [26]).
The classes of any semi-simple cohomological field theory belong to the tautological ring.
2.6. The trivial CohFT
An important case when the Dubrovin-Zhang hierarchy and the DR/DZ equivalence are fully understood is the case of the trivial CohFT [5, 6]. In particular, this implies the following special cases of Theorem 2.3:
Lemma 2.6.
The following three statements hold:
  1. (1)
    For any $(g,n,m)$ such that $g\geq 0$, $n\geq 1$, $m\geq 2$, and for any monomial $\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ we have
    (2.13) $\deg\left(\int_{{\overline{\mathcal{M}}}_{g,n+m}}\left(\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}\right)B^{m}_{g,n}\right)\leq 2g-2+m.$
  2. (2)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-1+n>$ 0, we have
    (2.14) $\deg\left(\int_{{\overline{\mathcal{M}}}_{g,n+1}}(B^{1}_{g,n}-A^{1}_{g,n})\right)\leq 2g-1.$
  3. (3)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-2+n>$ 0 we have
    (2.15) $\deg\left(\int_{{\overline{\mathcal{M}}}_{g,n}}(B^{0}_{g,n}-\frac{1}{\sum_{i=1}^{n}a_{i}}\pi_{*}A^{1}_{g,n})\right)\leq 2g-2.$
Proof.
The second and the third statement are literally equivalent to the statements of the strong DR/DZ correspondence in this case, see [5, 6, 16]. The first statement is equivalent by [16, Section 4] to the fact that the second and higher derivatives of the potential $F(t_{0},t_{1},\dots)$ corresponding to the string solution are known to be differential polynomials in the dependent variables in this case, see e. g. [18]. ∎
Remark 2.7.
Note that in the case $m=1$ we don’t have a ready statement on the degree of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+1}}\psi_{n+1}^{b}(B^{1}_{g,n}-A^{1}_{g,n})$ for $b>0$. The reason for this is that the first derivatives of the potential $F$ with respect to the variables $t_{b}$, $b>0$, play no direct role in the construction of integrable systems in this case.
3. The master relation in the Gorenstein quotient
3.1. Main definitions and statements
The definition and overall idea to use what we call master relation comes from a parallel paper [4], where we use a close relative of this relation to prove the conjectures on the so-called $\Omega$-classes posed in [3].
3.1.1. Pre-stable star rooted trees
Fix $m\geq 1$, $n\geq 1$, and $g\geq 0$. Let $\mathrm{PSSRT}_{g,n,m}$ be the set of pre-stable star rooted trees, that is, the rooted trees with one root vertex, where the frozen legs $\sigma_{n+1},\dots,\sigma_{n+m}$ are attached, and all other vertices are connected by an edge to the root (hence the term “star”). We also demand that no regular legs are attached to the root vertex.
The graphs that we obtain are quite similar to the graphs that form a subset of $\mathrm{SRT}_{g,n,m}$, and we extend the definitions applying the same notation for various concepts related to these graphs. However, since we allow non-stable vertices, $\mathrm{PSSRT}_{g,n,m}$ is not quite a subset of $\mathrm{SRT}_{g,n,m}$.
3.1.2. Classes assigned to vertices
Let $T\in\mathrm{PSSRT}_{g,n,m}$. As in the case of trees in $\mathrm{SRT}_{g,n,m}$, we assign to each $v\in V(T)$ the moduli space of curves ¯${\overline{\mathcal{M}}}_{g(v),|H(v)|}$, where the first $|H_{+}(v)|$ marked points correspond to the positive half-edges attached to $v$ and ordered in an arbitrary but fixed way and the the last $|H_{-}(v)|$ marked points correspond to the negative half-edges attached to $v$, also ordered in some arbitrary but fixed way. The cases of non-stable pairs $(g(v),|H(v)|)$ will be treated separately, but informally one can think of a natural extension of the classes we use for stable vertices to the unstable moduli spaces.
For the root vertex, we consider the class
(3.1) $\Psi(v_{r})$ $\coloneqq\prod_{i=1}^{|E(T)|}\frac{1}{1-a(h_{i})\psi_{i}}\in R^{*}({\overline{\mathcal{M}}}_{g(v_{r}),|E(T)|+m})\otimes_{\mathbb{Q}}Q$
Note that for pre-stable star rooted trees, $H_{+}(v_{r})$ can be identified with $E(T)$. In the exceptional unstable case $g(v_{r})=0,m=1,|E(T)|=1$ we formally assign to the root vertex the following class:
(3.2) $a(h_{1})^{-1}\in R^{-1}({\overline{\mathcal{M}}}_{0,2})\otimes_{\mathbb{Q}}Q[a(h_{1})^{-1}],$
where the negative cohomological degree is just formally assigned to allow to treat this case non-exceptionally in what follows.
For a non-root root vertex $v$, we consider the class
(3.3) ${\mathbb{D}}(v)$ $\coloneqq\frac{\lambda_{g(v)}\mathrm{DR}_{g(v)}\big{(}a(h_{1}),\dots,a(h_{|H_{+}(v)|}),-a(v)\big{)}}{1-a(v)\psi_{|H(v)|}}\in R^{*}({\overline{\mathcal{M}}}_{g(v),|H(v)|})\otimes_{\mathbb{Q}}Q$
Note that in this case $a(v)=-a(h_{H(v)})=a(h_{1})+\cdots+a(h_{|H_{+}(v)|})$. In the exceptional unstable case $g(v)=0,H_{+}(v)=1$ we formally assign to this vertex the following class:
(3.4) $a(h_{1})^{-1}\in R^{-1}({\overline{\mathcal{M}}}_{0,2})\otimes_{\mathbb{Q}}Q[a(h_{1})^{-1}],$
where, as it was for the root vertex, the negative cohomological degree is just formally assigned to allow to treat this case non-exceptionally in what follows.
3.1.3. Classes assigned to trees
Now, let us assume that $02g-2+n+m>$ 0 and assign to a tree $T\in\mathrm{PSSRT}_{g,n,m}$ a class $\Xi(T)\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q[u,u^{-1}]$. Here $u$ is a new formal variable to control the dimension. We have:
(3.5) $\Xi(T)\coloneqq u^{2g-2+m}\left(\prod_{e\in E(T)}\frac{a(e)}{u}\right)(b_{T})_{*}\left(\left(\sum_{d=-1}^{\infty}\frac{\Psi(v_{r})_{d}}{(-u)^{d}}\right)\otimes\bigotimes_{v\in V\setminus\{v_{r}\}}\left(\sum_{d=-1}^{\infty}\frac{{\mathbb{D}}(v)_{d}}{u^{d}}\right)\right),$
where $(b_{T})_{*}$ is the boundary pushforward map from $\bigotimes_{v\in V(T)}R^{*}({\overline{\mathcal{M}}}_{g(v),|H(v)|})$ to $R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ extended by linearity to the rational functions in $a_{i}$ and Laurent polynomials in $u$. In the case of pre-stable trees $b_{T}$ is assumed to contract the unstable components. Note that under the assumption $02g-2+n+m>$ 0 the dependence of the resulting formula on $a_{i}$’s is purely polynomial, so $\Xi(T)\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q[u,u^{-1}]$.
Remark 3.1.
Note that $[u^{2g-2+m-d}]\Xi(T)\in R^{d}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$ and $\Xi(T)_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
Remark 3.2.
As a side remark, note that in the case $02g-2+n+m=$ 0, that is, $g=0,n=m=1$, we have just one tree with two vertices, both unstable, and we can extend the definition of the class $\Xi(T)$ given above to produce the class $a_{1}^{-1}u^{0}\in R^{-1}({\overline{\mathcal{M}}}_{0,2})\otimes_{\mathbb{Q}}Q[a_{1}^{-1}][u,u^{-1}]$.
3.1.4. Classes assigned to sets of trees
Let $02g-2+n+m>$ 0. Consider the following Laurent polynomial in a formal variable $u$ and a polynomial in the variables $a_{1},\dots,a_{n}$:
(3.6) $\Xi^{m}_{g,n}\coloneqq\sum_{T\in\mathrm{PSSRT}_{g,n,m}}\Xi(T).$
Remark 3.3.
Note that $(\Xi^{m}_{g,n})_{d}\coloneqq[u^{2g-2+m-d}]\Xi^{m}_{g,n}\in R^{d}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q$ and $(\Xi^{m}_{g,n})_{d}$ is a homogeneous polynomial of degree $d$ in $a_{1},\dots,a_{n}$.
Conjecture 3.4 (master relation).
For any $g\geq 0$, $m,n\geq 1$, $02g-2+n+m>$ 0,
(3.7) $\Xi^{m}_{g,n}\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})\otimes_{\mathbb{Q}}Q[u].$
In other words, for any $d\geq 2g-1+m$ we have
(3.8) $(\Xi^{m}_{g,n})_{d}=0.$
The master relation is equivalent to Conjecture 2.1, the precise statement of this equivalence and its proof are delayed to the next section, see Theorem 4.5. For our applications, that is, for the case of tautological CohFTs and F-CohFTs, in particular for all semi-simple CohFTs, a weaker statement is sufficient:
Theorem 3.5 (master relation in the Gorenstein quotient).
For any $g\geq 0$, $m,n\geq 1$, $02g-2+n+m>$ 0, and for any $\alpha\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ we have
(3.9) ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha\Xi^{m}_{g,n}\in Q[u].$
In other words, for any $\alpha\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})$ we have
(3.10) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha\Xi^{m}_{g,n}|_{u=1}\leq 2g-2+m.$
The latter statement can also be rephrased as the vanishing of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha(\Xi^{m}_{g,n})_{d}$ for $d\geq 2g-1+m$ and any $\alpha\in R^{*}({\overline{\mathcal{M}}}_{g,n+m})$, and we call this vanishing the master relation in the Gorenstein quotient.
The rest of this section is devoted to a reduction of a proof of this theorem to the following statement:
Lemma 3.6.
Theorem 3.5 holds for any $\alpha$ if it holds for the classes $\alpha=\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}$ for any $b_{1},\dots,b_{m}\in\mathbb{Z}_{\geq 0}$.
3.2. Intersection with a divisor
For practical computations it is convenient to extend the definition above to $n=0$, where $\Xi^{m}_{g,0}$ is defined for $02g-2+m>$ 0 and is set to zero.
Let $\rho_{1}\colon{\overline{\mathcal{M}}}_{g-1,n+m+2}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the boundary map.
Lemma 3.7.
We have
(3.11) $\rho_{1}^{*}(\Xi^{m}_{g,n})_{d}=(\Xi^{m+2}_{g-1,n})_{d}.$
Remark 3.8.
Note that if $d\geq 2g-1+m$, then $d\geq 2(g-1)-1+(m+2)$, so the degree remains in the range that should vanish in the Gorenstein quotient once it was there before the pull-back.
Let $\rho_{2}\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+m_{1}+1}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+m_{2}+1}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the boundary map and assume that $m_{1}$ and $m_{2}$ are non-zero.
Lemma 3.9.
We have
(3.12) $\rho_{2}^{*}(\Xi^{m}_{g,n})_{d}=\sum_{d_{1}+d_{2}=d}(\Xi^{m_{1}+1}_{g_{1},n_{1}})_{d_{1}}\otimes(\Xi^{m_{2}+1}_{g_{2},n_{2}})_{d_{2}}.$
In this formula we assume that the arguments of $\Xi$-classes correspond to the marked points on the corresponding components.
Remark 3.10.
Note that if $d\geq 2g-1+m$, then either $d_{1}\geq 2g_{1}-1+(m_{1}+1)$ or $d_{2}\geq 2g_{2}-1+(m_{2}+1)$, so the degree on at least one of the two components is in the range that should vanish in the Gorenstein quotient once the degree of the original class was in the vanishing range before the pull-back.
Let $\rho_{3}\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+m+1}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+1}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the boundary map and assume that all points with the labels ${n+1},\dots,{n+m}$ are on the first component, and $n_{2}$ points on the second component have labels in $I\subset\{1,\dots,n\}$, $|I|=n_{2}$.
Lemma 3.11.
We have
(3.13) $\rho_{3}^{*}(\Xi^{m}_{g,n})_{d}$ $=\sum_{d_{1}+d_{2}=d}(\Xi^{m+1}_{g_{1},n_{1}})_{d_{1}}\otimes(\Xi^{1}_{g_{2},n_{2}})_{d_{2}}$
$\quad+(\Xi^{m}_{g_{1},n_{1}+1})_{d-2g_{2}}(\{a_{j}\}_{j\not\in I},a_{I})\otimes\lambda_{g_{2}}\mathrm{DR}_{g_{2}}\Big{(}\{a_{i}\}_{i\in I},-a_{I}\Big{)}.$
In this formula we assume that the arguments of $\Xi$-classes correspond to the marked points on the corresponding components. In the last summand we use $a_{i}\coloneqq\sum\nolimits_{i\in I}a_{i}$.
Remark 3.12.
Note that if $d\geq 2g-1+m$, then in the first summand either $d_{1}\geq 2g_{1}-1+(m+1)$ or $d_{2}\geq 2g_{2}-1+1$, so the degree on at least one of the two components is in the range that should vanish in the Gorenstein quotient once the degree of the original class was in the vanishing range before the pull-back. In the second summand, $d-2g_{2}\geq 2g_{1}-1+m$ once $d\geq 2g-1+m$.
The proof of all three lemmas is a straightforward application of the instructions on intersection of tautological classes given in [20] and a formula for the intersection of a double ramification cycle with a divisor in [17].
3.3. Intersection with a kappa class
Lemmata in Section 3.2 reduce the computation of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\alpha\Xi^{m}_{g,n}$ for an arbitrary class $\alpha$ to the computation of the same intersection numbers for the classes $\alpha=\kappa_{c_{1},\dots,c_{l}}\prod_{i=1}^{n+m}\psi_{i}^{b_{i}}$. To this end, we have the following statement:
Lemma 3.13.
We have
(3.14) ¯$\int_{{\overline{\mathcal{M}}}_{g,n+m}}\kappa_{c_{1},\dots,c_{l}}\prod_{i=1}^{n+m}\psi_{i}^{b_{i}}\Xi^{m}_{g,n}=\int_{{\overline{\mathcal{M}}}_{g,n+l+m}}\prod_{i=1}^{n}\psi_{i}^{b_{i}}\prod_{i=1}^{l}\psi_{n+i}^{c_{i}+1}\prod_{i=1}^{m}\psi_{n+l+i}^{b_{n+i}}\Xi^{m}_{g,n+l}|_{a_{n+1}=\cdots=a_{n+l}=0}.$
Proof.
Let $\pi_{n+1}\colon{\overline{\mathcal{M}}}_{g,n+1+m}\to{\overline{\mathcal{M}}}_{g,n+m}$ be the map that forgets the $(n+1)$-st marked point and shifts the labels of the last $m$ marked points by $-1$. Then $\pi_{n+1}^{*}\Xi^{m}_{g,n}=\Xi^{m}_{g,n+1}|_{a_{n+1}=0}$. In order to obtain (3.14) we iterate this formula $l$ times and use the projection formula for $\kappa_{c_{1},\dots,c_{l}}=(\pi_{n+1}\cdots\pi_{n+l})_{*}\prod_{i=1}^{l}\psi_{n+i}^{c_{i}+1}$. ∎
Remark 3.14.
Equation  (3.14) preserves the degree, that is, it is bounded by $2g-2+m$ on the left hand side if and only if it is bounded by $2g-2+m$ on the right hand side. In particular, together with Lemmata of Section 3.2, this reduces Theorem 3.5 to a weaker statement, where the class $\alpha$ is a product of $\psi$-classes.
3.4. Intersection with a $\psi$ class at a regular leg
Lemma 3.15.
For any $i=1,\dots,n$, we have:
(3.15) ${a_{i}\psi_{i}}(\Xi^{m}_{g,n})_{d}=(\sigma_{i})^{*}(\Xi^{m+1}_{g,n-1}(\{a_{j}\}_{j\not=i}))_{d+1}-(\Xi^{m}_{g,n})_{d+1}$
$+\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ n_{1}+n_{2}=n\\ i\in I\subseteq\{1,\dots,n\}\\ |I|=n_{2}\end{subarray}}a_{I}(\rho_{3})_{*}\left((\Xi^{m}_{g_{1},n_{1}+1})_{d-2g_{2}}(\{a_{j}\}_{j\not\in I},a_{I})\otimes\lambda_{g_{2}}\mathrm{DR}_{g_{2}}\Big{(}\{a_{i}\}_{i\in I},-a_{I}\Big{)}\right).$
Here in the first term by $\sigma_{i}$ we mean the relabeling of the points
(3.16) $(1,\dots,n+m)\to(1,\dots,i-1,i+1,\dots,n+m,i).$
In the last sum we use $a_{I}\coloneqq\sum\nolimits_{j\in I}a_{j}$ and assume that $02g_{1}-1+n_{1}+m>$ 0 and $02g_{2}-1+n_{2}>$ 0, $\rho_{3}$ is the boundary map described above.
Proof.
We just apply the following formula adapted from [17] to each term on the left hand side where the $i$-th marked point lies on a double ramification cycle, and then regroup the terms:
(3.17) $(a_{i}\psi_{i}+a\psi_{n+1})\lambda_{g}\mathrm{DR}_{g}(a_{1},\dots,a_{n},-a)=$
$\sum_{\begin{subarray}{c}g_{1}+g_{2}=g\\ n_{1}+n_{2}=n\\ i\in I\subseteq\{1,\dots,n\}\\ |I|=n_{2}\end{subarray}}a_{I}\rho_{*}\left(\lambda_{g_{1}}\mathrm{DR}_{g_{1}}(\{a_{j}\}_{j\not\in I},-a,a_{I})\otimes\lambda_{g_{2}}\mathrm{DR}_{g_{2}}(\{a_{i}\}_{i\in I},-a_{I})\right),$
where $a=\sum_{i=1}^{n}a_{i}$, $a_{I}=\sum_{i\in I}a_{i}$, and $\rho\colon{\overline{\mathcal{M}}}_{g_{1},n_{1}+2}\times{\overline{\mathcal{M}}}_{g_{2},n_{2}+1}$ is the boundary map that glues the last marked points on both components into a node and maps the first $(n_{1}+1)$ points on the first component to the points with the indices $j\not\in I$ and $(n+1)$ and the first $n_{2}$ points on the second component to $i\in I$. ∎
Remark 3.16.
Note that once on the left hand side of (3.15) $d\geq 2g-1+m$, then in all three summands on the right hand side of Equation (3.15) we have $d+1\geq 2g-1+(m+1)$, $d+1\geq 2g-1+m$, and $d-2g_{2}\geq 2g_{1}-1+m$ in the first, second, and third summand respectively.
Lemma 3.15 together with Remark 3.16 imply that the desired vanishing in the intersection numbers on the right hand side of Equation 3.14 follows from the vanishing of
(3.18) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+m}}\prod_{i=1}^{m}\psi_{n+i}^{b_{1}}\Xi^{m}_{g,n}|_{u=1}$
in degrees $\geq 2g-1+m$ for all $g\geq 0$, $n,m\geq 1$ and all $b_{1},\dots,b_{m}\in\mathbb{Z}_{\geq 0}$. Together with Lemma 3.13 and Lemmata of Section 3.2, this completes the proof of Lemma 3.6.
4. Proof of the main theorem
4.1. Strategy of proof of Theorem 2.3
We first give a short outline of the proof and then we expand it to make it more precise.
4.1.1. Step 1
Using a combinatorial argument of [4] adapted to our case, we state a lemma which controls the difference between either: the $B$-class and the $\Xi$-class when $m\geq 2$, or the $A-B$-class and the $\Xi$-class when $m=1$.
4.1.2. Step 2
A first consequence of this lemma is that the main theorem (Theorem 1.2) is equivalent to the master relation in the Gorenstein quotient (Theorem 3.5).
4.1.3. Step 3
We prove the master relation in the Gorenstein quotient. By a sequence of explicit intersections with tautological classes, we reduced Theorem 3.5 for any $\alpha$ to $\alpha=\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}$, see Lemma 3.6. To prove this special case, we use that the main theorem is already established in some special cases, see Lemma 2.6, together with a second use of the lemma of step $1$.
4.2. Step $1$: the key combinatorial lemma
We describe a result in [4] in quite general terms adapted to the needs of our argument.
Lemma 4.1 (Corollary of [4]).
The following two statements hold:
  1. (1)
    Let $m\geq 2$, $g\geq 0$, $n\geq 1$. Then, for any $d\geq 2g-1+m$,
    (4.1) $(B^{m}_{g,n})_{d}-(-1)^{d}(\Xi^{m}_{g,n})_{d}$
    is represented as a linear combination of tautological classes supported on graphs in $\mathrm{SRT}_{g,n,m}$ with at least one edge, where in each summand either one non-root vertex $v$ is decorated by $(\Xi^{1}_{g(v),|H_{+}(v)|})_{d(v)}$ with $d(v)\geq 2g(v)$ or the root vertex is decorated by $(B^{m}_{g(v_{r}),|H_{+}(v_{r})|})_{d(v_{r})}$ with $d(v_{r})\geq 2g(v_{r})-1+m$.
  2. (2)
    Let $m=1$, $g\geq 0$, $n\geq 1$, $02g-1+n>$ 0. Then, for any $d\geq 2g$,
    (4.2) $(B^{1}_{g,n}-A^{1}_{g,n})_{d}-(-1)^{d}(\Xi^{1}_{g,n})_{d}$
    is represented as a linear combination of tautological classes supported on graphs in $\mathrm{SRT}_{g,n,1}$ with at least one edge, where in each summand either one non-root vertex $v$ is decorated by $(\Xi^{1}_{g(v),|H_{+}(v)|})_{d(v)}$ with $d(v)\geq 2g(v)$ or the root vertex is decorated by $(B^{1}_{g(v_{r}),|H_{+}(v_{r})|}-A^{1}_{g(v_{r}),|H_{+}(v_{r})|})_{d(v_{r})}$ with $d(v_{r})\geq 2g(v_{r})$.
Remark 4.2.
In all cases, the structure of the tautological classes supported on the trees $T$ in $\mathrm{SRT}_{g,n,m}$ used in Lemma 4.1 is similar to the one used in the definition of $B^{m}_{g,n}$: each vertex is decorated by a class in $R^{*}({\overline{\mathcal{M}}}_{g(v),|H(v)|})\otimes_{\mathbb{Q}}Q$, whose component of cohomological degree $d$ is a homogeneous polynomial in $a_{i}$’s of degree $d$ (this can be $\Xi^{1}_{g(v),|H_{+}(v)|}$, $\Psi(v)$, or ${\mathbb{D}}(v)$, or eventually $(B^{m}_{g(v_{r}),|H_{+}(v_{r})|}-\delta_{m,1}A^{1}_{g(v_{r}),|H_{+}(v_{r})|})$ for the root vertex), and we multiply the push-forward $(b_{T})_{*}$ of the tensor product of such classes by $\prod_{e\in E(T)}a(e)$.
Remark 4.3.
As we see from the statement of Lemma 4.1, is it more convenient to use the class $\tilde{\Xi}^{m}_{g,n}\coloneqq\sum_{d\geq 0}(-1^{d})(\Xi^{m}_{g,n})_{d}$.
Remark 4.4.
Since the vertices of the graphs in $\mathrm{SRT}_{g,n,m}$ with at least one edge have strictly less negative Euler characteristic than $2g-2+m+n$, Lemma 4.1 allows to perform various inductive arguments. In particular, this lemma has two immediate corollaries, both proved by induction on $(g,n)$ with fixed $m$.
4.3. Step $2$: equivalence between the master relation and DR/DZ
It follows by induction using the key lemma that the master relation is equivalent to Conjecture 2.1, more precisely:
Theorem 4.5.
The following equivalences hold:
  1. (1)
    For any $(g,n)$ such that $g\geq 0$, $n\geq 1$, $02g-1+n>$ 0
    $\begin{cases}\deg\Xi_{g,n}^{1}\leq 2g-1,\\ \deg\Xi_{g^{\prime},n^{\prime}}^{1}\leq 2g^{\prime}-1,\ \text{for all}\ g^{\prime}\leq g,\,n^{\prime}\leq n,\,2g^{\prime}+n^{\prime}<2g+n\end{cases}\qquad\Longleftrightarrow\qquad\\ \begin{cases}\deg(B^{1}_{g,n}-A^{1}_{g,n})\leq 2g-1,\\ \deg(B^{1}_{g^{\prime},n^{\prime}}-A^{1}_{g^{\prime},n^{\prime}})\leq 2g^{\prime}-1,\ \text{for all}\ g^{\prime}\leq g,\,n^{\prime}\leq n,\,2g^{\prime}+n^{\prime}<2g+n.\end{cases}$
  2. (2)
    Fix $m\geq 2$. For any $g\geq 0$, $n\geq 1$, $02g-2+n+m>$ 0,
    $\begin{cases}\deg\Xi_{g,n}^{m}\leq 2g-2+m,\\ \deg\Xi_{g^{\prime},n^{\prime}}^{m}\leq 2g^{\prime}-2+m,\ \text{for all}\ g^{\prime}\leq g,\,n^{\prime}\leq n,\,2g^{\prime}+n^{\prime}<2g+n\\ \deg\Xi_{g^{\prime},n^{\prime}}^{1}\leq 2g^{\prime}-1,\ \text{for all}\ g^{\prime}\leq g,\,n^{\prime}\leq n,\,2g^{\prime}+n^{\prime}<2g+n\end{cases}\quad\Longleftrightarrow\quad\\ \begin{cases}\deg B^{m}_{g,n}\leq 2g-2+m,\\ \deg B^{m}_{g^{\prime},n^{\prime}}\leq 2g^{\prime}-2+m,\ \text{for all}\ g^{\prime}\leq g,\,n^{\prime}\leq n,\,2g^{\prime}+n^{\prime}<2g+n\\ \deg(B^{1}_{g^{\prime},n^{\prime}}-A^{1}_{g^{\prime},n^{\prime}})\leq 2g^{\prime}-1,\ \text{for all}\ g^{\prime}\leq g,\,n^{\prime}\leq n,\,2g^{\prime}+n^{\prime}<2g+n.\end{cases}$
In particular, this equivalence holds in the Gorenstein quotient. Thus it suffices to prove the master relation in the Gorenstein quotient, that is Theorem 3.5, to prove the main theorem.
4.4. Step 3: proof of the master relation in the Gorenstein quotient
We prove Theorem 3.5 for $\alpha=\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}$ which is implies Theorem 3.5 for any $\alpha$ by Lemma 3.6.
First, using the key lemma we deduce by induction the following statement:
Corollary 4.6.
The following statements hold
  1. (1)
    For any $g\geq 0$, $n\geq 1$, $02g-1+n>$ 0, and $b\geq 0$,
    (4.3) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+1}}\psi_{n+1}^{b}\left(B^{1}_{g,n}-A^{1}_{g,n}-\tilde{\Xi}^{1}_{g,n}\right)\leq 2g-1$
    if
    (4.4) $\deg\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+1}}\psi_{n+1}^{b}\left(B^{1}_{g^{\prime},n^{\prime}}-A^{1}_{g^{\prime},n^{\prime}}\right)\leq 2g-1$
    holds for any $g^{\prime}\geq 0$, $n^{\prime}\geq 1$, such that $g^{\prime}\leq g$, and $2g^{\prime}+n^{\prime}<2g+n$, and also
    (4.5) $\deg\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+1}}\tilde{\Xi}^{1}_{g^{\prime},n^{\prime}}\leq 2g^{\prime}-1$
    holds for any $g^{\prime}\geq 0$, $n^{\prime}\geq 1$, such that $g^{\prime}\leq g$, and $2g^{\prime}+n^{\prime}<2g+n$.
  2. (2)
    Fix $m\geq 2$. For any $g\geq 0$, $n\geq 1$, and $b_{1},\dots,b_{m}\in\mathbb{Z}_{\geq 0}$,
    (4.6) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+m}}\left(\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}\right)\left(B^{m}_{g,n}-\tilde{\Xi}^{m}_{g,n}\right)\leq 2g-2+m$
    if
    (4.7) $\deg\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+m}}\left(\prod_{i=1}^{m}\psi_{n^{\prime}+i}^{b_{i}}\right)B^{m}_{g^{\prime},n^{\prime}}\leq 2g^{\prime}-2+m$
    holds for any $g^{\prime}\geq 0$, $n^{\prime}\geq 1$, such that $g^{\prime}\leq g$, and $2g^{\prime}+n^{\prime}<2g+n$, and also
    (4.8) $\deg\int_{{\overline{\mathcal{M}}}_{g^{\prime},n^{\prime}+1}}\tilde{\Xi}^{1}_{g^{\prime},n^{\prime}}\leq 2g^{\prime}-1$
    holds for any $g^{\prime}\geq 0$, $n^{\prime}\geq 1$, such that $g^{\prime}\leq g$, and $2g^{\prime}+n^{\prime}<2g+n$.
We recall our convention that a polynomial with negative degree is the zero polynomial.
Then, we combine the statements of Lemma 2.6 and Corollary 4.6. Inductively, we obtain the following statement:
Lemma 4.7.
We have
(4.9) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+m}}\left(\prod_{i=1}^{m}\psi_{n+i}^{b_{i}}\right)\tilde{\Xi}^{m}_{g,n}\leq 2g-2+m$
for $m\geq 2$ and any $b_{1},\dots,b_{m}\in\mathbb{Z}_{\geq 0}$ and
(4.10) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+1}}\tilde{\Xi}^{1}_{g,n}\leq 2g-1.$
Note that we still don’t have the statement for the degree of ¯$\int_{{\overline{\mathcal{M}}}_{g,n+1}}\psi_{n+1}^{b}\tilde{\Xi}^{1}_{g,n}$ for any $b\in\mathbb{Z}_{\geq 0}$. However, by the pushforward formula which implies the string equation, we have
(4.11) ¯$\int_{{\overline{\mathcal{M}}}_{g,n+2}}\psi_{n+1}^{b}\psi_{n+2}^{0}\tilde{\Xi}^{2}_{g,n}=\int_{{\overline{\mathcal{M}}}_{g,n+1}}\psi_{n+1}^{b-1}\tilde{\Xi}^{1}_{g,n}+\left(\sum_{i=1}^{n}a_{i}\right)\int_{{\overline{\mathcal{M}}}_{g,n+1}}\psi_{n+1}^{b}\tilde{\Xi}^{1}_{g,n}.$
Hence, by induction on $b$, we obtain the following corollary of Lemma 4.7:
Corollary 4.8.
We have
(4.12) $\deg\int_{{\overline{\mathcal{M}}}_{g,n+1}}\psi_{n+1}^{b}\tilde{\Xi}^{1}_{g,n}\leq 2g-1$
for any $b\in\mathbb{Z}_{\geq 0}$.
This completes the proof of Theorem 3.5 by Lemma 3.6, and hence of Theorem 2.3.